The normal human body has an average core or interior temperature of 37C∘. At a skin temperature of 34 C∘, how much heat is removed through conduction? Assume a thermal conductivity of unperfused (i.e. bloodless) tissue of 18 (Cal x cm)/(m^2 x hr x C), 3 cm thickness of the tissue between the interior and the exterior of the body, and 1.5 m2 total surface area of the human body. 

Choices:

  • 36 Cal/hr
  • 29 Cal/hr
  • 27 Cal/hr
  • 18 Cal/hr
  • 22.5 Cal/hr
1
Expert's answer
2021-01-01T14:06:51-0500

"q=\frac{Q}{t}=\lambda\cdot\frac{\Delta T}{\Delta x}\cdot A=18\cdot\frac{37-34}{3}\cdot1.5=27(Cal/hr)" . Answer